zarkam21:

Besides protons, what other particles make up an atom? Write 2 - 3 sentences identifying the attractive forces that act between a proton and each of these other particles.

zarkam21:

Besides protons, neutrons and electrons make up an atom

Vocaloid:

good, what about the attractive forces between protons + electrons and protons + neutrons?

zarkam21:

The attractive forces between protons and electrons cancel out. The attractive forces between protons and neutrons attract?

Vocaloid:

it asks to identify what the forces are so what are the names of the forces between protons + neutrons and protons + electrons?

zarkam21:

The attractive forces between protons and electrons is called an electrostatic force The attractive forces between protons and neutrons is called a strong nuclear force.

Vocaloid:

good that's it

zarkam21:

Using a periodic table, identify the element whose atoms have 11 protons. Is it a metal, a nonmetal, or a metalloid? Write 1 - 2 sentences describing how an atom of this element changes when it forms an ionic bond.

Vocaloid:

any attempts so far?

Vocaloid:

ptable.com is a periodic table, find element 11 and identify whether it's a metal/nonmetal/metalloid based on its location, then that will tell you how it behaves during bonding

zarkam21:

The element whose atoms have 11 protons is a metal

Vocaloid:

good it would also be a good idea to write what the name of the element is

zarkam21:

The element whose atoms have 11 protons is named Sodium and it is a metal

zarkam21:

It is very reactive

Vocaloid:

good, and during ionic bond formation does it lose or gain electrons?

zarkam21:

lose

zarkam21:

actually gain

Vocaloid:

other way around, since it only has 1 e- in its outer shell it has to lose them so there you have it: sodium, metal, loses electrons during ionic bonding

zarkam21:

A robotic submersible is being brought back on board a ship. What happens to the buoyant force as the submersible is lifted out of the water? Write 2 - 3 sentences explaining your reasoning.

Vocaloid:

well, buoyant force is proportional to the amount of water being displaced by the object, when an object is being raised, water gets displaced below the object, thus exerting an upwards force on the object, raising the object

Vocaloid:

so the buoyant force should increase

zarkam21:

If 0.45 kg of iron at 140°C is added to 1.5 kg of water at 22°C, what will be the final temperature of the system?

Vocaloid:

q = mc delta T the iron is losing heat and the water is gaining heat therefore [m_iron][Ciron](Tf-Ti) = [m_water][Cwater](Tf-Ti) convert the temperatures to kelvin find the appropriate values of C for iron and water find the appropriate values of m for iron and water plug and solve for Tf

Vocaloid:

actually nvm about the Kelvin thing, you can find specific heat in terms of kg and degrees C

Vocaloid:

any progress?

zarkam21:

no :/

Vocaloid:

specific heat of iron = 450J/kg specific heat of water = 4186J/kg

Vocaloid:

plug in the mass of the iron the mass of the water the initial T of the water the initial T of the iron [m_iron][Ciron](Tf-Ti) = [m_water][Cwater](Tf-Ti) solve for Tf.

zarkam21:

[m_iron][Ciron](Tf-Ti) = [m_water][Cwater](Tf-Ti) 0.45*140=1.5*22

Vocaloid:

your equation is not complete. You are missing (Tf-Ti) on both sides. You also used the wrong value for Cwater.

zarkam21:

0.45*140(f-0.45)=1.5*22(f-1.5)

Vocaloid:

Ti is the initial temperature not the mass Also, Cwater is not 22

zarkam21:

0.45*140(f-140)=1.5*1.5(f-22)

zarkam21:

?

Vocaloid:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Vocaloid specific heat of iron = 450J/kg specific heat of water = 4186J/kg \(\color{#0cbb34}{\text{End of Quote}}\)

zarkam21:

0.45*140(f-140)=1.5*4186(f-22)

Vocaloid:

good forgot to mention that the iron is losing heat and the water is gaining heat so there needs to be a negative sign on the iron side -0.45*140(f-140)=1.5*4186(f-22) solve for f

zarkam21:

23.17

Vocaloid:

good (didn't get the notification for some reason ;_;)

zarkam21:

and it would be degrees c right?

Vocaloid:

yes